### v1

# Trust, but verify

Abstract
Decentralized AI agent networks, such as Gaia, allows individuals to run customized LLMs on their own computers and then provide services to the public. However, in order to maintain service quality, the network must verify that individual nodes are running their designated LLMs. In this paper, we demonstrate that in a cluster of mostly honest nodes, we can detect nodes that run unauthorized or incorrect LLM through social consensus of its peers. We will discuss the algorithm and experimental data from the Gaia network. We will also discuss the intersubjective validation system, implemented as an EigenLayer AVS to introduce financial incentives and penalties to encourage honest behavior from LLM nodes.
Keywords LLM  $\cdot$ open source  $\cdot$ decentralized inference  $\cdot$ cryptoeconomics
1 Introduction
Two years after OpenAI released ChatGPT, there is an explosion of open-source LLMs both in quality and quantity [Zhao et al. 2025]. The best open-source LLMs are now matching the best close-source LLMs in performance [Fourrier, Habib, Lozovskaya, Szafer & Wolf 2024], while offering much more variety in terms of model size, speed, and specific "skills", such as math and coding [Jiang, Wang, Shen, Kim & Kim 2024]. The rise of open source enables individual users to run customized LLMs and agent applications on their own computers without centralized gatekeepers. The decentralized LLM inference has many benefits, such as privacy, low cost, speed, and availability. But, more importantly, it fosters an ecosystem of customized LLM services based on models fine-tuned and supplemented by proprietary data and knowledge.
Decentralized AI inference networks, such as Gaia [Gaia Foundation 2024], allow individuals to pool together their computing resources to serve the in-demand customized LLMs to the public and share revenues with the model creators. Each Gaia node runs a set of AI models, including LLMs, and a knowledge base. It advertises and makes available its services to the public. However, as those networks must stay permissionless, to reduce censorship, protect privacy and lower the barrier of participation, it is possible for some Gaia nodes to cheat and run a different LLM or knowledge base than the ones it claims to run. A popular Gaia "domain", which is a sub-net of nodes that are supposed to run the same LLM and knowledge base, could have over 1,000 nodes at any given time. The network must have a mechanism to automatically detect and penalize the nodes that cheat.
2 Past research
Researchers have indicated that they could deterministically verify any computing machine using cryptographic algorithms.
Zero Knowledge Proof (ZKP) allows us to verify the outcome of a computation process without knowing the internals of the computing system [Goldwasser, Micali & Rackoff 1985]. In the context of LLM verification, we can build a ZKP circuit for an LLM, which represents the inference process. The LLM service provider (Gaia node) generates a proof for each incoming request (i.e., the LLM prompt) indicating that it has correctly executed the inference process [Sun, Li & Zhang 2024]. Any external party (validator) can then check the proof and confirm that it is generated by the specific LLM without knowing the exact software setup on the machine.
However, ZKP also has some severe limitations that make it impossible to use in production environments.
  • The ZKP circuit must be generated for each LLM. That requires vast engineering efforts since there are tens of thousands of open-source LLMs on Huggingface alone.
  • Even the state-of-the-art ZKP algorithms running on expensive Nvidia GPUs require 13 minutes to generate proof for a single inference execution for a small 13B LLM [Sun, Li & Zhang 2024]. That is 2 orders of magnitude slower (and 2 orders of magnitude more resource consuming) than the inference task itself. Since LLM inference is already compute-bound, spending 100×100$\times$100 × on compute just for verification is not feasible.
  • The memory cost of generating ZKP proofs is also staggering. A toy LLM with one million parameters would require over 25GB of RAM even with the most optimized algorithms [Ganescu & Passerat-Palmbach 2024]. Scaling that to LLMs of 10,000×10,000$\times$10 , 000 × of that size would require computing resources far beyond any reasonable computer.
  • If the LLM itself is open source, it is possible to fake the proof, which would render the ZKP entirely useless in most decentralized AI inference systems where open-source is a requirement.
A second cryptographic approach is to use hardware-based Trusted Execution Environment (TEE) that are built into CPUs and GPUs. TEE can generate signed hashes (i.e., attestation) for software and data on the processor. That ensures that the software and data exactly matches a specified version. Since the TEE is a hardware feature, it is impossible for the computer operator or owner to manipulate or fake a proof.
However, TEE too, is not ready for large-scale AI inference.
  • The TEE reduces raw hardware (CPU) performance by a factor of up to 2×2$\times$2 × [Dong & Wang 2025], which is very hard to accept since LLM inference is already a compute-bound task.
  • There are very few GPUs or AI accelerators that support TEEs. The high-end H-100 is the first Nvidia GPU with TEE support [Apsey, Rogers, O’Connor & Nertney 2023].
  • Even if we can verify that a specific LLM is running inside a TEE, it is still impossible to verify that an LLM server is actually using that LLM to serve Internet requests. There are many moving parts in an AI server that resides outside of the TEE.
  • Distributing private keys to decentralized and permissionless TEE devices is a major challenge that requires specialized software and operational procedures [Oasis 2025].
As we have seen above, the cryptographic approaches are slow, expensive, and generally impractical with today’s consumer-grade computer hardware. A more promising approach is to use cryptoeconomic mechanisms, which optimistically assume that most players in a decentralized inference network are honest and then use social consensus to detect dishonest players. Through staking and slashing, the network can incentivize honest players and punish dishonest players, and hence creating a virtuous cycle.
Since the LLMs are non-deterministic, we need a group of validators that continuously sample the LLM service providers, and vote on the answers to detect dishonest LLM service providers and also dishonest validators. The sampling, data collection and voting are all conducted off-chain through interactions between the validators and nodes. But the results are recorded on-chain and have cryptoeconomic consequences.
EigenLayer’s AVS [EigenLayer 2023] provides a set of smart contracts that allows independent operators and validators to stake crypto assets, and vote on intersubjective results from the sampling. Once a vote reaches consensus, known as an intersubjective consensus [EigenLayer 2024], the smart contracts automatically execute to reward or punish the operators and validators.
The voting rules and algorithms are the subject of this research. In the rest of this paper, we will first present hypotheses on how to distinguish dishonest LLM service providers (i.e., Gaia nodes) from a group of honest providers. We will then run experiments and validate the hypotheses through empirical data. We will also discuss AVS implementation design and considerations for such a system.
3 Hypotheses
Recall that a Gaia node is a network server that runs a set of AI models and knowledge bases. It promises to run the AI and knowledge base stack it advertises on the Gaia network.
We hypothesize that by asking questions to a set of Gaia nodes, and observing the statistical distributions of the answers, we will be able to detect outliers that run different LLMs or knowledge bases than the majority.
Specifically, when we ask a question to all nodes in a Gaia domain, the answers will form a tight cluster in a high-dimensional embedding space, and if there are outliers that lies far outside of the cluster, the corresponding nodes are likely running a different LLM or knowledge base than the ones required for this domain.
We formulate a set of questions $Q$ and a set of Gaia nodes $M$. In each experiment, we choose a question $q$ and send it to the Gaia node $m$ for an answer.
$q\in Q$ (1)
$m\in M$ (2)
The question is repeated $n$ times for each node $m$ to create a distribution of the answers. We will hence have $n$ answers. Each answer is passed to an LLM embedding model to generate a $z$ dimensional vector point that is representative of its semantic meaning. In the case of the gte-Qwen2-1.5B-instruct embedding model, $z$ is $1536$. The $z$ dimensional vector point from the $i$th answer of question $q$ on node $m$ is written as follows.
$X_{i}(q,m)=[x_{0,i}(q,m),...x_{z,i}(q,m)]$ (3)
We can compute the mean point of answers generated from question $q$ and node $m$ as follows.
$\overline{X}(q,m)=\frac{\left[\sum_{i=0}^{n-1}x_{0,i}(q,m),...\sum_{i=0}^{n-1}x_{z,i}(q,m)\right]}{n}$ (4)
For a given pair of $q,m$ and $q^{\prime},m^{\prime}$, the difference between the answers can be measured as the distance between the mean points.
$D(q,m,q^{\prime},m^{\prime})=\sqrt{\sum_{j=0}^{z-1}\left(\frac{\sum_{i=0}^{n-1}x_{j,i}(q,m)-\sum_{i=0}^{n-1}x_{j,i}(q^{\prime},m^{\prime})}{n}\right)^{2}}$ (5)
The consistency of the $n$ answers generated by a node $m$ for a question $q$ is measured as the standard deviation of the vector points.
$\sigma_{j}(q,m)=\sqrt{\frac{\sum_{i=0}^{n-1}\left(x_{j,i}-\frac{\sum_{i=0}^{n-1}x_{j,i}}{n}\right)^{2}}{n}}$ (6)
We can compute the Root-Mean-Square (RMS) to represent the internal variance of the distribution.
$\overline{\sigma}(q,m)=\sqrt{\frac{1}{z}\sum_{j=0}^{z-1}\sigma_{j}^{2}(q,m)}$ (7)
Our hypotheses are therefore as follows.
Hypothesis 1: When two Gaia nodes with different LLM or knowledge base configurations $m$ and $m^{\prime}$ answer the same question $q$, the answers form distinct distributions that are specific to the Gaia node. This allows an outside observer to easily tell which Gaia node generated a given answer.
$D(q,m,q,m^{\prime})>3\times(\overline{\sigma}(q,m)+\overline{\sigma}(q,m^{\prime}))$ (8)
Hypothesis 2: When a Gaia node $m$ answers two independent questions $q$ and $q^{\prime}$, the answers form distinct distributions that are specific to the question. This allows an outside observer to easily tell which question resulted a given answer.
$D(q,m,q^{\prime},m)>3\times(\overline{\sigma}(q,m)+\overline{\sigma}(q^{\prime},m))$ (9)
4 Experiments
A Gaia node has two key components that affect the responses that can be observed from an outside validator: the LLM or the knowledge base (vector database).
Hence we ran two sets of experiments to test the above hypotheses on LLM and knowledge base variations respectively. The source code and data to replicate all the analyses and results in this paper are available as an open source GitHub repo [Gaia Foundation 2025a].
4.1 Responses from different models
In the first experiment, we aim to test whether we can reliably distinguish responses from different LLMs. We set up 3 Gaia nodes for the following 3 open source LLMs.
  • Llama 3.1 8b by Meta AI [Grattafiori et al. 2024].
  • Gemma 2 9b by Google [Mesnard et al. 2024].
  • Gemma 2 27b by Google [Mesnard et al. 2024].
Each model was queried with 20 factual questions covering topics such as science, history, and geography (see Appendix A). Each question was repeated 25 times per model, yielding 500 responses per model and 1,500 responses in total.
4.2 Responses from different knowledge bases
In the second experiment, we test whether we can reliably distinguish responses from different knowledge bases used by the same LLM.
We set up two Gaia nodes with the same LLM, Gemma-2-9b by Google [Mesnard et al. 2024]. The knowledge bases are vector databases created from text files [Gaia Foundation 2025b]. In this experiment, we adopted knowledge bases from Paris and London Wikipedia pages respectively [Wikipedia 2025a, 2025b]. The knowledge bases are created with the nomic-embed-text-1.5 embedding model [Nussbaum, Morris, Duderstadt, and Mulyar, 2025]. The vector database files are available as follows.
  • Paris [Yuan 2024a].
  • London [Yuan 2024b].
Each knowledge node was queried with 20 factual questions with even coverage on Paris and London topics (see Appendix B). Each question was repeated 25 times per topic (or knowledge base), yielding 500 responses per knowledge base and 1,000 responses in total.
5 Results
For each response from the experiments, we did the following.
  • Used a standard system prompt: "You are a helpful assistant."
  • Recorded the complete text response
  • Generated an embedding vector using the gte-Qwen2-1.5B-instruct embedding model [Yang et al. 2024]
We then analyzed the responses in the embedding vector space.
5.1 Distinguishing different models
Figure 1 shows the internal consistency of answers from the same question for each LLM, measured by the RMS scatter.
Figure 1: Internal consistency of different LLMs measured by RMS scatter.
The consistency metrics reveal that Gemma-2-27b demonstrates the highest consistency in responses, with an RMS scatter of $0.0043$. Llama-3.1-8b shows the highest variation with an RMS scatter of $0.0062$.
To evaluate whether different models give different answers for the same question, Figure 2 shows the mean distance of answer clusters generated by pairs of LLMs.
Figure 2: Average distance vs RMS scatter for each question and model pair.
The distance matrix reveals that inter-model distances vary substantially across questions. For instance, the highest distance observed was $0.5291$ between Llama-3.1-8b and Gemma-2-27b for question Q13 ("Who wrote ’Romeo and Juliet’?"). The lowest non-zero distance was $0.0669$ between Gemma models for question Q14 ("What is the atomic number of oxygen?").
Critically, Table 1 shows that the distances between model pairs are $32-65$ times larger than the highest RMS scatter observed within any model ($0.0062$ for Llama-3.1-8b). This large separation between inter-model distances and intra-model variation indicates that Gaia nodes with different LLMs produce reliably distinguishable outputs, making them identifiable through their response patterns.
Table 1: Average distance $D_{ave}$ over max RMS scatter $\sigma_{max}$ for each pair of LLMs
Model pair $D_{ave}$ $D_{ave}/\sigma_{max}$
Gemma9b - Gemma27b $0.1558$ 32.5×32.5$\times$32.5 ×
Llama8b - Gemma9b $0.3129$ 65.2×65.2$\times$65.2 ×
Llama8b - Gemma27b $0.3141$ 65.4×65.4$\times$65.4 ×
5.2 Distinguishing different knowledge bases
Figure 3 shows the internal consistency of answers from the same question for each knowledge base, measured by the RMS scatter. The Paris knowledge base is denoted as kb_a, and the London knowledge base is denoted as kb_b in the figures.
Figure 3: Internal consistency of different knowledge bases measured by RMS scatter.
The consistency metrics show that the two knowledge bases generate responses of similar consistency. To evaluate whether nodes with different knowledge bases give different answers for the same question, Figure 4 shows the mean distance of answer clusters generated by the two knowledge bases.
Figure 4: Average distance vs RMS scatter for each question and knowledge base pair.
The distance matrix reveals that inter-knowledge base distances vary substantially across questions. For instance, the highest distance observed was $0.188$ between Paris and London knowledge bases for question Q18 ("How many bridges did Philip Augustus build in Paris in the late 12th century?"). The lowest non-zero distance was $0.037$ between knowledge bases for question Q19 ("What percentage of Paris’s salaried employees work in hotels and restaurants according to the document?").
Critically, the distances between knowledge base pairs (average $0.0862$) are about $5-26$ times larger than the RMS scatter observed within a knowledge base ($0.0072$). This large separation between inter-knowledge base distances and intra-knowledge base variation indicates that Gaia nodes with different knowledge bases produce reliably distinguishable outputs, making them identifiable through their response patterns.
Our hypotheses are validated by the experimental results.
6 Discussions
The ratios between inter-model / inter-knowledge base distances and intra-model / intra-knowledge base variations are statistically substantial. Further research would be needed to determine how this ratio might be affected by factors such as changing hardware, varying load conditions, or minor model updates. These initial results suggest a potential basis for statistical differentiation between models and knowledge bases.
The two Gemma models are more similar to each other (distance of $0.1558$) than either is to Llama (distances around $0.31$). While still 32×32$\times$32 × greater than the observed internal variation, this suggests that distinguishing between models within the same family might present different challenges than distinguishing between models from different families.
Paradoxically, different knowledge bases produce more similar answers (distances around $0.1$) than different LLM models. The reason is probably due to that our two knowledge bases are both about European capitals. While the answers about London or Paris are meaningfully different, they might actually look similar to the embedding models. That said, the inter-knowledge base distances are still substantially larger than scatters within a knowledge base, allowing observers to distinguish between knowledge bases from statistical properties of the answers.
Finally, different questions demonstrated varying levels of differentiation between models or knowledge bases. For example, Q13 ("Who wrote ’Romeo and Juliet’?") showed the greatest distance between models ($0.5291$ between Llama 8b and Gemma 27b). Q18 ("How many bridges did Philip Augustus build in Paris in the late 12th century?") showed the greatest distance between knowledge bases ($0.188$ between London and Paris). Those variations in the effectiveness of test questions are important in practical applications.
7 System design and implementation
Gaia domains are permissionless collections of AI inference nodes (Gaia node). Each Gaia node in a Gaia domain agrees to run the exact LLM (GGUF model file) and knowledge base (Qdrant vector snapshot) specified by the domain owner. That allows the Gaia domain, as a whole, to provide consistent services to the public.
In this paper, we propose an AVS design that is tailored to the Gaia network architecture. The AVS is organized into multiple "operator sets".
  • The operator set 0 is a special group for AVS validators. An AVS validator is responsible for polling Gaia nodes with questions and performing statistical analysis to detect outliers within each Gaia domain. In order to join operator set 0, the validator must be approved by the Gaia DAO organization.
  • The operator sets 1 to $n$ are mapped to Gaia domains. All Gaia nodes in each domain form a single operator set.
The AVS operates on Epochs. Each epoch lasts 12 hours. During the epoch, each validator would perform the following tasks repeatedly.
  • Poll each Gaia node in a domain using a random question from a question set that is specially designed for this domain. The question set will be provided by the domain owner.
  • Record all responses, time-outs, and error messages.
  • Detect outlier nodes based on responses.
  • Create time-encrypted and signed messages for the results.
  • Post the results on EigenDA [EigenDA 2025].
At the end of each Epoch, a randomly chosen validator will act as the aggregator and perform a map-reduce function on all messages posted by all validators in that Epoch. The result is a set of flags for each Gaia node, such as
  • outlier – The Gaia node produces responses that are statistical outliers compared to other nodes in the same domain, indicating that it might be running the incorrect LLM or knowledge base required for the domain.
  • slow – The Gaia node responds much slower (i.e., outside of three standard deviations from the mean response time) than nodes in the same domain.
  • timeout – The Gaia node timed out in one or more requests.
  • error 500 – The Gaia node returned HTTP error code 500, indicating an internal server error, in one or more requests.
  • error 404 – The Gaia node returned HTTP error code 404, indicating unavailable resources, in one or more requests.
  • error other – The Gaia node returned any other HTTP error code in one or more requests.
The cumulative status of a Gaia node (i.e., AVS operator) at the end of each Epoch determines its AVS award or punishment from this Epoch. For example, if a Gaia node maintains a good status, without any flags, in the past 10 epochs, it will receive the regular AVS award. If it is flagged as an outlier or an error 500 in the current epoch, or flagged as slow for several epochs in a row, it might be suspended from both the AVS award and domain participation for the next several epochs. For malicious actors, the AVS might even slash their stakes. The detailed rules on awards and punishments will be determined by the Gaia network.
Furthermore, participating domains could also ask the AVS validators to onboard new Gaia nodes. Currently, a Gaia node can apply to join a domain. The domain owner reviews and approves the application. The AVS validators could take over this role. They could poll the candidate node with requests specific to the domain. Gaia nodes that conform to the domain requirements for LLM, knowledge base, and response speed will be automatically admitted as the validator results are posted in EigenDA.
8 Conclusions
In this paper, we demonstrated that statistical analysis of LLM outputs could reliably signal the model and knowledge base in use. That allows decentralized LLM inference networks to use EigenLayer AVS to verify LLM outputs intersubjectively and detect outliers as potential bad actors. Combined with cryptoeconomic incentives and penalties in AVSes, it is possible to scale this LLM verification approach to large networks.